OpenStudy (rock_mit182):

n IS A positive integer, a perfect number is equal to the sum of its proper positive divisors, that is, the sum of its positive divisors excluding the number itself. if p and 2^p-1 are primes, then a perfect number is ?

OpenStudy (rock_mit182):

the expression is : \[2^{p} -1\]

OpenStudy (tkhunny):

I'm not certain we have a question, here. If p is prime, does it have any "Proper Divisors"?

OpenStudy (rock_mit182):

yeah but how am i going to find the expression for a perfect number given p and\[2^{p}-1\]

OpenStudy (rock_mit182):

i guess since they're prime numbers the only divisor they colud have is one, but how i make the proof and find the expression for the perfect number ?

OpenStudy (tkhunny):

Like I said, I'm not sure we have a question.

OpenStudy (tkhunny):

Oh, we get to count "1" in the proper divisors. We just don't get to count n.

OpenStudy (rock_mit182):

for example if i multiply \[p * 2^{p} +1\] that would be a perfect number ?

OpenStudy (rock_mit182):

if so how am i going to proof that

OpenStudy (tkhunny):

Note: It wants "A" perfect number, not "THE" perfect number. You had better not multiply anything like that. Remember your Order of Operations. Try \(p\cdot \left(2^{p} - 1\right)\) It's proper divisors are what?

OpenStudy (tkhunny):

Answer my question. What are the proper divisors of that product?

OpenStudy (rock_mit182):

i mean i saw the possible answer in wikipedia: "Euclid proved that\[2^{p-1}(2^{p}-1)\]

OpenStudy (rock_mit182):

but i don't understand how to get there from my question

OpenStudy (tkhunny):

Why are you not answering my question? \(p\cdot (2^{p} - 1)\) has proper divisors \(1,\;p,\;and\;2^{p}-1\) Do you believe?

OpenStudy (rock_mit182):

yes i know, but since has to to be a perfect number, couldn't include p and 2^p-1

OpenStudy (rock_mit182):

@ranga @mathmale @whpalmer4

OpenStudy (rock_mit182):

@wio

OpenStudy (tkhunny):

You're not seeing it. Outside the box. This course will require nonstandard thinking from your brain. You have to see outside where you have been, before. \(p\cdot\left(2^{p} - 1\right)\) is a lovely number. Let's just talk about this number for a moment. Let's call this number "Steve". Steve has Proper Divisors of \(1,\;p,\;and\;2^{p}-1\) If Steve is equal to the sum of Steve's proper divisors, then Steve is a Perfect Number. Can we track down a specific value for Steve? \(Steve\;=\;p\cdot\left(2^{p}-1\right)\;=\;1 + p + \left(2^{p}-1\right)\) Can we do anything with that? Steve is lonely. He's having an identity crisis.

OpenStudy (whpalmer4):

It's as simple as 1, 2, 3...